OpenStudy (anonymous):

A flower bed is in the shape of a triangle with one side twice the length of the shortest side, and the third side is 15 feet more than the length of the shortest side. Find the dimensions if the perimeter is 131.

OpenStudy (academicgurusinc):

Hello xtriickyy Answer: side one = 29 side two = 2 * 29 side three = 15+ 29 Your question will be featured in our Tutoring Thursday YouTube session. We will send you the link once it's completed so that you will have a visual of the process. Regards, AcademicGurusInc. Twitter (@Academic_Gurus) Facebook (AcademicGurusInc)

OpenStudy (academicgurusinc):

Hello The video has been uploaded. https://www.youtube.com/watch?v=YPEXs_CLHc0 Regards, AcademicGurusInc.